Shift report 20030508 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu May 08 2003 - 22:07:56 EDT

  • Next message: Apache: "Shift report 20030509 08:00-16:00"
    ------------------------------------------------------------
    Date: Thu May  8 22:07:56 2003, Shift: 16:00-24:00
    Supervisor: rk, with: 
    Log book#14 pages 101 - 101
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam until 18.40, then quite nice store, which ended at 21.50. However there were beam studies from 20.00 which affected the collision rates...
    
    2. Done before beam studies:
    Run#   tr.2   tr.3   tr.6
    8867     .2k   9k      .9k   (tr.6 scaled down by 3)
    8868    1.5k  62k    15k
    With beam studies underway:
    Run#   tr.2   tr.3   tr.6
    8869     .7k  32k     8k
    8870   the rates were very bad - I finished the run right         .      away
    8871     .4k  19k     5k
    
    3. No detector problems observed
    
    4. The shift ended earlier - there should be only 6 bunches. FV said it's OK to end shift after 8pm
    
    5. TPCs and DCs are in the park mode, magnets are off.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu May 08 2003 - 22:08:32 EDT